OpenStudy (anonymous):

What is the factored form of 8x2 + 3x − 5?

OpenStudy (anonymous):

8x2 + 3x − 5=8x2 + 8x-5x − 5=8x(x+1)-5(x+1)=(x+1)(8x-5)